OpenStudy (anonymous):

If f and g are continuous functions such that then in terms of g is:

OpenStudy (anonymous):

g'(x)=f(4x)

OpenStudy (anonymous):

no i mean g'(x)=4f(x)

OpenStudy (anonymous):

then \[\int\limits_{1}^{2}f(2x)dx\] in terms of g is

OpenStudy (agent0smith):

So there's no confusion... If f and g are continuous functions such that g'(x)=4f(x), then \[\Large \int\limits\limits_{1}^{2}f(2x)dx \] in terms of g is?

OpenStudy (anonymous):

correct

OpenStudy (agent0smith):

g'(x)=4f(x) Since the derivative of g is 4f, then the integral of 4f(x) must be g(x) (i pulled out the constant 4) \[\Large \int\limits g'(x) dx= 4 \int\limits f(x) dx\] or using limits \[\Large 4 \int\limits_{a}^{b} f(x) dx = g(b) - g(a)\] divide both sides by 4 \[\Large \int\limits\limits_{a}^{b} f(x) dx =\frac{ 1 }{4 } \left[ g(b) - g(a) \right]\]

OpenStudy (agent0smith):

Now... lets make a substitution first. u = 2x so du = 2 dx or du/2 = dx Change the limits using u = 2x. When x=2, u = 4. When x=1, u = 2. \[\Large \frac{ 1}{ 2}\int\limits\limits\limits\limits\limits\limits\limits_{2}^{4}f(u)du = \frac{ 1 }{ 2 } \times \frac{ 1 }{4 }\left[ g(4) - g(2) \right]\]